OpenStudy (anonymous):

An investment of P dollars made at the simple interest rate r for t years will grow to a total T = P + Prt. For an investment of 10,000 to grow to at least 12,000 in two years, what is the minimum interest rate at which it can be invested?

OpenStudy (theeric):

Hi! A lot of math speak in this one. Does it have you confused?

OpenStudy (anonymous):

yea, sorry for bugging you though. It's from my math homework x__x

OpenStudy (anonymous):

I got lost when it asked for the minimum rate

OpenStudy (theeric):

It's not a problem :) Well, your total investment, \(T\), will be \(\sf \color{blue}{at\ least}\) \(12,000\). It can be \(12,000\) \(\sf\color{blue}{or\ greater}\). So, in math speak, \(T\ge12,000\). Now, you want to solve for the interest rate, \(r\), and see what is the smallest value that you can have for it (that is the minimum).

OpenStudy (anonymous):

ahh I see

OpenStudy (theeric):

:) So, are you ready to tackle the rest of the problem on your own? I'll still be here.

OpenStudy (anonymous):

I'll try to

OpenStudy (anonymous):

so the equation should start out like T = 10,000 + 12,000(r)2 right?

OpenStudy (theeric):

Close! You have \(P\) as 10,000 in one place and 12,000 in another!

OpenStudy (theeric):

I think that 12,000 should be a 10,000, since that's what \(P\) is. The 12,000 is \(T\).

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

so now it's 12,000 = 10,000 + 2r? just to be sure

OpenStudy (theeric):

12,000 = 10,000 + (10,000)2r :)

OpenStudy (anonymous):

ahh ok thanks ^^

OpenStudy (theeric):

:)

OpenStudy (anonymous):

alright thanks for the help ^^

OpenStudy (theeric):

You're welcome! Did you solve for \(r\)?

OpenStudy (anonymous):

yea , but I'm kind of unsure of it though ^^|| I got 0.1 = r. I'm sure if I got it right or not

OpenStudy (anonymous):

~not sure~

OpenStudy (theeric):

Yeah! \(.1\ge r\), so the smallest \(r\), a.k.a. the "minimum interest rate," is \(.1\). Sometimes rates are percentages, so you might want to say \(.1=10\%=r_\text{minimum}\)

OpenStudy (theeric):

So I agree with you!:)

OpenStudy (anonymous):

oh ok thank you :3

OpenStudy (theeric):

No problem.